OpenStudy (patd):

Solve: -2(2x + 5) - 3 = -3(x - 1)

OpenStudy (jiteshmeghwal9):

take all the variable terms on left hand side & all constant terms on right hand side

OpenStudy (patd):

I tried to combine like terms by multiplying the parenthesis on the left side by -2, but after that I have no idea what to do.

OpenStudy (jiteshmeghwal9):

k! wht do u gt after multiplying the parentheses by -2 on left hand side ?

OpenStudy (patd):

I got -4x+(-10)-3=-3x-3 when I combined like terms for both sides.

OpenStudy (jiteshmeghwal9):

okay collecting the variable terms on left hand side -4x+3x=-3+3+10

OpenStudy (patd):

So would the answer be x=-10 ?